3) Find the inverse Fourier transform of X(w) = jw +3 -w2 + j3w + 2 jw +3 (jw + 1) (w + 2)2 X(W) =
Answer #1

X(W) = jw + 3 -w f 3 je ta x(W) = jw+ 3 j w + 3 (w) + 3 ja ta ( w)taj w t jw +2 ý 6+3 jwt3 jul je taj +1(Juta) (swt) (Jw+2) DX ) jW+3 (w+1) (Juraj Do Padbal Fachons jw + 3 jwta ( w ta (wtil (jwtag ACJW tas + BCJ W H) (jwta) + cj wri) = jw +3 jw =-2 (Proof Tousier transform of ē a uct) = X(+) uc)-S1;120 Eco Xcwl J MCA) e-swt... dt e ult) el jot (1) -jot 8no - Cat jcolt -cat